OpenStudy (anonymous):

Solve for w: (w-3/5)+(3w-8/4)=5

OpenStudy (anonymous):

:) sure

OpenStudy (anonymous):

\[\frac{ w-3 }{5 }+\frac{ 3w-8 }{ 4 }=5\]

OpenStudy (anonymous):

good. so do u know what to do 1st?

OpenStudy (anonymous):

Find the common multiple of 5 and 4?

OpenStudy (anonymous):

yup. so what is it?

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

good so then..

OpenStudy (anonymous):

Multiply each set by 20?

OpenStudy (anonymous):

no. hold up just one sec

OpenStudy (anonymous):

so u know the part that has a denominator of 5?

OpenStudy (anonymous):

you'd multiply the numerator by 4 so its be 4w-12

OpenStudy (anonymous):

& the part where it has a denominator of 4? youd multiply it denominator by 5 to get 15w-25

OpenStudy (anonymous):

\[\frac{ 4w-3 }{ 5 }+\frac{ 15w-25 }{ 4 }\]

OpenStudy (anonymous):

Is it supposed to look like this?

OpenStudy (anonymous):

why would it be 4w-3???

OpenStudy (anonymous):

itd be 4w-12 right/? cuz 4*3=12...

OpenStudy (anonymous):

& actuallythe other side would be 15w-40

OpenStudy (anonymous):

oops okay

OpenStudy (anonymous):

what about the denominators?

OpenStudy (anonymous):

yea. umm lemme see

OpenStudy (anonymous):

sorry had to help someone quickly

OpenStudy (anonymous):

so here it should look like this: